OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = -6 and a6 = -6,144? (1 points) Answer -327,680 -393,216 -458,752 -524,288

OpenStudy (anonymous):

a_n*r^n = nth term

OpenStudy (anonymous):

ok please teach me how i set it up?

OpenStudy (anonymous):

what do i put in for "n"?

OpenStudy (anonymous):

hold on i am on it.

OpenStudy (anonymous):

formula is off a bit formula is this a*r^(n-1) = nth term so first term is -6 cuzz if we plug it in a * r^( 1-1) = -6 ^ that is 0 a * r^0 =-6 ^ that is 1 cuzz anynumber to the 0th is 1 just a rule a=-6 so now we know a and plug it into the a6 6th term a *r^(n-1) = nth term -6 *r^(6-1) = -6144 -6 * r^5 = 6144 sfr r = -4 so now we have all variables a = - 6 and r is 4 so plug it into formula a*r^(n-1) = nth term -6*4^(9-1)= -393216

hartnn (hartnn):

the problem was that you solved it entirely by yourself @timo86m instead of making her solve...

OpenStudy (anonymous):

no offense @timo86m but he's right... I don't learn that way

OpenStudy (anonymous):

The problem is that someone might not always be here for you so you need the answer for confirmation.